OpenStudy (anonymous):

12+2e^(2x)=19 help please

OpenStudy (anonymous):

an example ques :3e-4/7 +2/3 = 2e+3/3 ans:3e-2e=3/3+4/7-2/3 e=(21+12-14)/21 e=19/21

OpenStudy (anonymous):

@Rohangrr your example has nothing to do with her question! \[\LARGE 12+2e^{2x}=19\] \[\LARGE 2e^{2x}=19-12\] \[\LARGE e^{2x}=\frac72\] \[\LARGE \ln(e^{2x})=\ln\left(\frac72\right) \] \[\LARGE 2x\cdot \underbrace{\ln(e)}_{\text{This is}\; 1}=\ln\left(\frac72\right) \] \[\LARGE x=\frac{\ln\left(\frac72\right)}{2} \]

OpenStudy (anonymous):

that's an ugly answer anyway... lol

OpenStudy (anonymous):

@Kreshnik it is just AN EXAMPLE!

OpenStudy (anonymous):

@Rohangrr I'm sure she'd do YOUR example easily! ... I assume she was confused about "e" and "x" ... anyway, examples are good. Medal goes to you. ;)